OpenStudy (anonymous):

✔️21y (5)✔️49y HELP PLEASE !!

OpenStudy (the_fizicx99):

.-. What...?

OpenStudy (anonymous):

Sqrt25y(5)sqrt49y

OpenStudy (anonymous):

mathway.com

OpenStudy (the_fizicx99):

Yeah I still don't know what that is .-. Try making it more clear?

OpenStudy (the_fizicx99):

\(\ \sf \sqrt{25y} \times 5\sqrt{49y} \) ?